f dustless chalk Instituted a quality control program A manufacturer o monitor chalk density. The sample standard deviations of densities for 24 different subgroups, each consisting of n 8 chalk specimens, were as follows: This data has been coded so that you may copy and paste it into R with the name k.sdevs. k.sdevs c(0.207, 0.313, 0.097, 0.186, 0.233, 0.209, 0.319, 0.290, 0.143 0.212, 0.054, 0.146, 0.274, 0.348, 0.161, 0.216, 0.090, 0.056) 0.152, 0.231, 0.274, 0.121, 0.385, 0.188. mean(k.sdevs) # Construct your own Control Chart. LCL 0.15 #### change this value to be the same as your calculateed LCL UCL 0.35 #### change this value to be the same as your calculated UCL plot(k.sdevs) (k evs)) sd eal abline(h UCL, Ity 2) Calculate limits for an S chart. (Round your answers four decimal places.) an- LCL = UCL-
Construct the chart UCL 0.4 0.4 UCL 0.3 0.3 0.2 0.2 "LCL 0.1 0.1 LCL Subgroup 0.0 Subgroup 0.0 5 10 15 20 25 5 10 15 20 25 -0.1 -0.1 S UCL 0.4 0.4 UCL 0.3 0.3 0.2 0.2 LCL 0.1 0.1 P LCL - Subgroup 25 0.0 15 20 5 10 Subgroup 0.0 15 25 10 20 -0.1 Check for out-of-control points. Most points are between these limits, so the process appears to be in control with respect to variability. All points are between these limits, so the process appears to be out of control with respect to variability. OMost points are between these limits, so the process appears to be out of control with respect to variability. The value of s on the 17th sample lies above the UCL, so the process appears to be out of control. All points are between these limits, so the process appears to be in control with respect to variability.
Answer #1

4-895-0.20 3.25; 24 24 So, cental imit s0.204 co LcL 3-33-a,2 an Heae n8 LaL 0.20-3.204-a 0.204 1-C0.965 20 4-3 o ESR .965 7.ES; = 4.895-0.388 = 4.507 3a4.507-0 196 93 LCL =0.196-3 0.19 -Co.965) 2 O.965 0.196-0.16 F0.036 UCL 0.196+o.16 0.356 04t VCL